OpenStudy (anonymous):

The implicit form of a linear equation is: y=8-1/2x What is the explicit form?

OpenStudy (amistre64):

what defines implicit and explicit formulas?

OpenStudy (anonymous):

Well, implicit is when the equation is like, jumbled up. Explicit is in the right format of y=mx+c.

OpenStudy (amistre64):

your function doesnt seem jumbled up to me tho ...

OpenStudy (amistre64):

ax+by =c, is an implicit linear format y=-a/b x +c/b, explicitly defines y in terms of x

OpenStudy (anonymous):

I get that bit ^-^ So would it be y=1/2x+8 ?